Use the given data to find the equation of the regression line. Examine the scatterplot and identify a characteristic of thea. Using the pairs of values for all 10 points, find the equation of the regression line. b. After removing the point with cob. What is the equation of the regression line for the set of ​points?

The data show the chest size and weight of several bears. Find the regression equation, letting chest size be the independentThe best predicted weight for a bear with a chest size of 48 inches is .......nothing pounds.

The data show the bug chirps per minute at different temperatures. Find the regression equation, letting the first variable bThe best predicted temperature when a bug is chirping at 3000 chirps per minute is .........F.
Answer #1

1)

ΣX ΣY Σ(x-x̅)² Σ(y-ȳ)² Σ(x-x̅)(y-ȳ)
total sum 110.00 135.32 110.00 137.0 110.00
mean 10.00 12.30 SSxx SSyy SSxy

Sample size,   n =   11      
here, x̅ = Σx / n=   10.000          
ȳ = Σy/n =   12.302          
SSxx =    Σ(x-x̅)² =    110.0000      
SSxy=   Σ(x-x̅)(y-ȳ) =   110.0      
              
estimated slope , ß1 = SSxy/SSxx =   110/110=   1.0000      
intercept,ß0 = y̅-ß1* x̄ =   12.3018- (1 )*10=   2.3018      
              
Regression line is, Ŷ=   2.30 + (   1.00 )*x
==============================

2)

using all ten points

a)

ΣX ΣY Σ(x-x̅)² Σ(y-ȳ)² Σ(x-x̅)(y-ȳ)
total sum 56.00 44.00 20.40 20.4 -14.40
mean 5.60 4.40 SSxx SSyy SSxy

Sample size,   n =   10      
here, x̅ = Σx / n=   5.600          
ȳ = Σy/n =   4.400          
SSxx =    Σ(x-x̅)² =    20.4000      
SSxy=   Σ(x-x̅)(y-ȳ) =   -14.4      
              
estimated slope , ß1 = SSxy/SSxx =   -14.4/20.4=   -0.7059      
intercept,ß0 = y̅-ß1* x̄ =   4.4- (-0.7059 )*5.6=   8.3529      
              
Regression line is, Ŷ=   8.353   + (   -0.706   )*x

b) after removing point (2,8)

Regression line is, Ŷ=   4.000   + (   0.000   )*x
=======================

3)

ΣX ΣY Σ(x-x̅)² Σ(y-ȳ)² Σ(x-x̅)(y-ȳ)
total sum 311.00 1724.00 222.83 24591.3 2282.33
mean 51.83 287.33 SSxx SSyy SSxy

Sample size,   n =   6      
here, x̅ = Σx / n=   51.833          
ȳ = Σy/n =   287.333          
SSxx =    Σ(x-x̅)² =    222.8333      
SSxy=   Σ(x-x̅)(y-ȳ) =   2282.3      
              
estimated slope , ß1 = SSxy/SSxx =   2282.3333/222.8333=   10.2423      
intercept,ß0 = y̅-ß1* x̄ =   287.3333- (10.2423 )*51.8333=   -243.5610      
              
Regression line is, Ŷ=   -243.6 + (   10.2 )*x

Predicted Y at X=   48   is          
Ŷ=   -243.56096   +   10.24233   *48=   248.1

========================

4)

ΣX ΣY Σ(x-x̅)² Σ(y-ȳ)² Σ(x-x̅)(y-ȳ)
total sum 5615.00 462.90 132398.83 671.4 9082.75
mean 935.83 77.15 SSxx SSyy SSxy

Sample size,   n =   6      
here, x̅ = Σx / n=   935.833          
ȳ = Σy/n =   77.150          
SSxx =    Σ(x-x̅)² =    132398.8333      
SSxy=   Σ(x-x̅)(y-ȳ) =   9082.8      
              
estimated slope , ß1 = SSxy/SSxx =   9082.75/132398.8333=   0.0686      
intercept,ß0 = y̅-ß1* x̄ =   77.15- (0.0686 )*935.8333=   12.9505      
              
Regression line is, Ŷ=   12.95 + (   0.0686 )*x